OpenStudy (anonymous):

ues substition method to fin solution sets: I need step by step help 2a+4b=-24 a=20-2b

OpenStudy (mathstudent55):

In the substitution method, you first need to solve one equation for one variable. In this case, that step is already done because the second equation shows what a is equal to. In other words, the second equation is already soved for a.

OpenStudy (mathstudent55):

The second step is to substitute the variable a in the first equation with what a is equal to in the second equation. Since a = 20 - 2b, everywhere you see a in the firtst equation, wrtite 20 - 2b

OpenStudy (anonymous):

i have a+b=-12 a+2b=20 is this correct?

OpenStudy (mathstudent55):

No. The first equation is 2a + 4b = -24 Where you see a, write (20 - 2b) You should have no a's left there.

OpenStudy (anonymous):

i get -4 is this correct?

OpenStudy (anonymous):

as my finale answer

OpenStudy (anonymous):

for b

OpenStudy (mathstudent55):

You ask for step by step help. I give you step by step help, but you skip a bunch of steps. You still haven't shown what you got after the substitution.

OpenStudy (mathstudent55):

The first equation is 2a + 4b = -24 Where you see a, write (20 - 2b)

OpenStudy (anonymous):

i did

OpenStudy (mathstudent55):

Where is it then? It should look like this: 2a + 4b = -24 2(20 - 2b) + 4b = -24